# Closing a catenary loop: the lariat chain, the string shooter, and the heavy elastica

### v1

@@ 6.1 Catenary loops
where we have used the fact that $d_{s}(\sigma-v^{2})=d_{s}\sigma$. The two fields $\sigma$ and $\theta$ depend on $s$, while the parameters $v$ and $D$ are uniform. A space-fixed boundary or patching point is non-material, so a jump condition [27] obtains there:
@@
Solutions with drag can be obtained following the derivation in [3], as detailed in A.1 of the present paper. The tension is eliminated to form a single equation for $\theta$ with one parameter $D$, which is then integrated to obtain the results:
@@
which may be compared with [3, equations 30-31]. The integration constant $C$ relates the total length and subtended angle of a segment of curve. These solutions, whose singularities at the vertical depend on the size of $D$, are valid as written for ranges of $\theta$ such that the parenthetical quantity is positive. We will typically be working with clockwise-parameterized upper branches, where $\theta$ starts near $\frac{\pi}{2}$ and decreases to near $\frac{\pi}{2}$, and $C$ is negative; lower branches are constructed from these by symmetry operations. This and other issues related to numerical implementation and patching of upper and lower branch solutions to form a closed loop are detailed in A.2-A.3.
Solutions of the shifted tension (8) and curvature (7) equations respectively undergo distinct bifurcations at $D=1$ and $D=2$, related to the nature of singularities at vertical orientation. [^7] Some of the behavior of solutions can also be inferred from an expansion of (4-5) around the vertical $\theta=-\frac{\pi}{2}$ if the arc length is finite there, requiring $D>1$. Denoting the deviation in angle and arc length from this point respectively as $\Theta$ and $S$, one obtains an approximate dynamical system $d_{S}\Theta=\Theta/(D-1)S$ with solutions $\Theta\sim S^{1/(D-1)}$, the latter also obtained by [24]. For $1<D<2$, the exponent is greater than unity, and both $\Theta\rightarrow 0$ and $d_{S}\Theta\rightarrow 0$ as $S\rightarrow 0$, with non-unique solutions at $S=0$, a feature missed in [3]. For $D>2$, the exponent is less than unity but still positive, and $\Theta\rightarrow 0$ and $d_{S}\Theta\rightarrow\infty$ as $S\rightarrow 0$. This analysis suggests the presence of an infinite sequence of bifurcations through which each higher derivative of $\theta$ transitions from vanishing to blowing up at the vertical; for example, the second and first derivatives of curvature $d_{S}^{3}\Theta$ and $d_{S}^{2}\Theta$ should respectively undergo transitions at $D=4/3$ and $D=3/2$. The tendency of a quantity to vanish or blow up may become weak near bifurcations, as reflected in the relevant exponents. Referring to the regimes discussed in Section 2,
@@
As also discussed in [3], special cases occur at the two bifurcations. When drag and weight are equal ($D=1$), the shifted tension takes on a finite value at the vertical, and when drag is twice the weight ($D=2$), the curvature takes on a finite value at the vertical. The physical meaning of these special cases remains unclear, and neither has been experimentally realized.
@@ 6.2 Regularization by bending stiffness
It is easy enough to physically hang a static loop from a support, or launch a slow-moving one, yet the physics of a catenary forbids such solutions. The most straightforward resolution of this issue comes from the addition of a bending stiffness, which allows for inflectional and vertical points. Such points are clearly present in experiments, including for example in [24, figure 1c] or [22, figure 2a], the latter reminiscent of [28, figure 4d] and [6, plate III figure 3]. The resulting problem is that of the heavy elastica or stiff catenary [29]. In the static limit, the physical situation is that of holding an elastic loop at one point, as between two fingers that supply a force and moment. We will also still allow for the presence of a kink.
@@
is potentially redundant, as it provides the same information as the normal projection of the linear momentum balance. However, we use it in our numerical formulation (detailed in A.4), and we will refer to it later in a discussion of global balances (Section 6.3). Defining the quantity appearing in the derivative in (9) as $\bm{f}\equiv\bm{n}-v^{2}\bm{\hat{e}}_{t}$, the shifted tension is now $\bm{f}\cdot\bm{\hat{e}}_{t}=\sigma-v^{2}+E(d_{s}\theta)^{2}$. Jump conditions [27] for these equilibrium configurations are ${\bf{R}}+\llbracket\bm{f}\rrbracket=\bm{0}$ and ${\bf{M}}+\llbracket\bm{m}\rrbracket=\bm{0}$, where the latter condition, allowing for the presence of a singular source of angular momentum $\bf{M}$} at the support, is identical to a static boundary condition for this beam model without rotary inertia. The only jump is at the support point, as there is no longer a need for patching at the vertical. It is apparent from from (10) that the $d_{s}^{3}\theta$ term, the second derivative of curvature, is what will balance the product of shifted tension and cuvature at the vertical, where $\cos\theta=0$.
@@ 6.3 Global balances
the latter also following directly from the definition of $\bm{m}$. These two quantities can be identified as $\bf{R}$} and $\bf{M}$} at the support point, where a jump $\llbracket q\rrbracket=q(0)-q(1)$. The force is purely vertical, balancing gravity; the drag does not contribute. In the string limit ($E=0$), it is not possible to apply a moment, and as pointed out in several previous works [4, 22, 24], the integral moments about the support due to drag and gravity balance each other; we can also infer that the moment about the center of mass due to drag balances that due to the support force.
It is illustrative to integrate the tangential projection of the linear momentum balance for axially moving equilibria (9), not simplified into (11), but rather rearranged into the form of a pseudomomentum balance [30, 27, 31] (see also [32, 33] for strings):
@@ 7 Open questions
While a small amount of bending offers the simplest physically plausible regularization of the vertical singularity, the resulting numerical problem is quite stiff, making comparison with experiment difficult. This suggests the development of a matched asymptotic analysis, likely to be quite different from prior analyses for heavy elastica in which the boundary layer is due to clamping [34, 35, 36]. We expect the “dolphin nose” [23], visible in experiments, to emerge at lower bending stiffnesses than those examined here.
@@ 8 Summary and Conclusions
The effects of drag on string shooter solutions can be briefly summarized as follows. For a stationary shape, drag forces do not contribute to the normal force balance, but they can change the equilibrium shape and tension. At low values, they do not change the qualitative features of the catenary, in particular its asymptotic approach to verticality, and the blowup of tension of opposite signs on upper and lower branches. Additional physics, such as bending elasticity, is necessary to construct loop solutions. At moderate values, the inertially-shifted tension vanishes along with the curvature at verticality at finite arc length, allowing patching of loop solutions. At high values, the curvature blows up at verticality, with patching still possible.
We have corrected several errors in recent works on axially moving string loops in the presence of gravity. These entail misunderstandings of the role of drag forces, bifurcations, the nature of the vertical singularity, and admissibility of certain solutions. Several example solutions were constructed, either by implementing previously derived analytical results [4, 3], or numerically adding bending stiffness. We have also attempted to place the problem in context alongside several other fundamental problems in flexible structure mechanics.
@@ A.1 Derivation of catenary solutions
where $C$ is a constant of integration. For general values of $D$, these expressions are valid as written for the range $\frac{\pi}{2}<\theta<\frac{\pi}{2}$, corresponding in our work to an upper branch of clockwise-travelling string. The compact expressions in the middle are preferred, but the rewritten expressions on the right are intended for further comparison with the work of Gregory [4], and were obtained using the identities $\sec\theta+\tan\theta=\tan\left(\frac{\pi}{4}+\frac{\theta}{2}\right)$ and $\sec\theta-\tan\theta=\cot\left(\frac{\pi}{4}+\frac{\theta}{2}\right)$.
@@ A.2 Patching of catenary solutions
It turns out that the above procedure for matching positions only works if the upper and lower portions of the curve are the same length, $\ell_{\text{up}}=\ell_{\text{down}}=\tfrac{\ell}{2}=\tfrac{1}{2}$, between the launch point and the vertical singularity. This result is, surprisingly, independent of the launch and return angles, so that, for example, the lengths of any red and green curves in the family shown in [24, figure 3] are the same. This equal length condition was found analytically by Gregory [4]. We reproduce an equivalent demonstration in the following subsection.
@@ A.3 Gregory’s demonstration of equal upper and lower lengths
Borrowing the shorthand notation of [4], we define the quantity appearing in equations (17)-(21) as $\tau\equiv\tan\left(\frac{\pi}{4}+\frac{\theta}{2}\right)$, and note that $\tau(-\tfrac{\pi}{2})=0$. Using (19-21) with a launch angle such that $\frac{\pi}{2}\leq\theta_{1}=\theta_{\text{launch}}<\frac{\pi}{2}$ and $\frac{\pi}{2}\leq\theta_{2}\leq\theta_{\text{launch}}$, we may obtain the shape of the upper portion of the loop ($s_{\text{up}}$, $x_{\text{up}}$, $y_{\text{up}}$) and find $C_{\text{up}}$ by inverting (19) with a chosen $\theta_{\text{launch}}$ and $\ell_{\text{up}}=s(-\frac{\pi}{2})$. Explicitly, we have

### v2

@@ 6.1 Catenary loops
where we have used the fact that $d_{s}(\sigma-v^{2})=d_{s}\sigma$. The two fields $\sigma$ and $\theta$ depend on $s$, while the parameters $v$ and $D$ are uniform. A space-fixed boundary or patching point is non-material, so a jump condition OReilly17book obtains there:
@@
Solutions with drag can be obtained following the derivation in ChakrabartiHanna16 , as detailed in A.1 of the present paper. The tension is eliminated to form a single equation for $\theta$ with one parameter $D$, which is then integrated to obtain the results:
@@
which may be compared with [ChakrabartiHanna16, , equations 30-31]. The integration constant $C$ relates the total length and subtended angle of a segment of curve. These solutions, whose singularities at the vertical depend on the size of $D$, are valid as written for ranges of $\theta$ such that the parenthetical quantity is positive. We will typically be working with clockwise-parameterized upper branches, where $\theta$ starts near $\frac{\pi}{2}$ and decreases to near $\frac{\pi}{2}$, and $C$ is negative; lower branches are constructed from these by symmetry operations. This and other issues related to numerical implementation and patching of upper and lower branch solutions to form a closed loop are detailed in A.2-A.3.
Solutions of the shifted tension (8) and curvature (7) equations respectively undergo distinct bifurcations at $D=1$ and $D=2$, related to the nature of singularities at vertical orientation. [^7] Some of the behavior of solutions can also be inferred from an expansion of (4-5) around the vertical $\theta=-\frac{\pi}{2}$ if the arc length is finite there, requiring $D>1$. Denoting the deviation in angle and arc length from this point respectively as $\Theta$ and $S$, one obtains an approximate dynamical system $d_{S}\Theta=\Theta/(D-1)S$ with solutions $\Theta\sim S^{1/(D-1)}$, the latter also obtained by Daerr19 . For $1<D<2$, the exponent is greater than unity, and both $\Theta\rightarrow 0$ and $d_{S}\Theta\rightarrow 0$ as $S\rightarrow 0$, with non-unique solutions at $S=0$, a feature missed in ChakrabartiHanna16 . For $D>2$, the exponent is less than unity but still positive, and $\Theta\rightarrow 0$ and $d_{S}\Theta\rightarrow\infty$ as $S\rightarrow 0$. This analysis suggests the presence of an infinite sequence of bifurcations through which each higher derivative of $\theta$ transitions from vanishing to blowing up at the vertical; for example, the second and first derivatives of curvature $d_{S}^{3}\Theta$ and $d_{S}^{2}\Theta$ should respectively undergo transitions at $D=4/3$ and $D=3/2$. The tendency of a quantity to vanish or blow up may become weak near bifurcations, as reflected in the relevant exponents. Referring to the regimes discussed in Section 2,
@@
As also discussed in ChakrabartiHanna16 , special cases occur at the two bifurcations. When drag and weight are equal ($D=1$), the shifted tension takes on a finite value at the vertical, and when drag is twice the weight ($D=2$), the curvature takes on a finite value at the vertical. The physical meaning of these special cases remains unclear, and neither has been experimentally realized.
@@ 6.2 Regularization by bending stiffness
It is easy enough to physically hang a static loop from a support, or launch a slow-moving one, yet the physics of a catenary forbids such solutions. The most straightforward resolution of this issue comes from the addition of a bending stiffness, which allows for inflectional and vertical points. Such points are clearly present in experiments, including for example in [Daerr19, , figure 1c] or [Taberlet19, , figure 2a], the latter reminiscent of [WatsonWang81, , figure 4d] and [Aitken1878, , plate III figure 3]. The resulting problem is that of the heavy elastica or stiff catenary Wang86 . In the static limit, the physical situation is that of holding an elastic loop at one point, as between two fingers that supply a force and moment. We will also still allow for the presence of a kink.
@@
is potentially redundant, as it provides the same information as the normal projection of the linear momentum balance. However, we use it in our numerical formulation (detailed in A.4), and we will refer to it later in a discussion of global balances (Section 6.3). Defining the quantity appearing in the derivative in (9) as $\bm{f}\equiv\bm{n}-v^{2}\bm{\hat{e}}_{t}$, the shifted tension is now $\bm{f}\cdot\bm{\hat{e}}_{t}=\sigma-v^{2}+E(d_{s}\theta)^{2}$. Jump conditions OReilly17book for these equilibrium configurations are ${\bf{R}}+\llbracket\bm{f}\rrbracket=\bm{0}$ and ${\bf{M}}+\llbracket\bm{m}\rrbracket=\bm{0}$, where the latter condition, allowing for the presence of a singular source of angular momentum $\bf{M}$} at the support, is identical to a static boundary condition for this beam model without rotary inertia. The only jump is at the support point, as there is no longer a need for patching at the vertical. It is apparent from from (10) that the $d_{s}^{3}\theta$ term, the second derivative of curvature, is what will balance the product of shifted tension and cuvature at the vertical, where $\cos\theta=0$.
@@ 6.3 Global balances
the latter also following directly from the definition of $\bm{m}$. These two quantities can be identified as $\bf{R}$} and $\bf{M}$} at the support point, where a jump $\llbracket q\rrbracket=q(0)-q(1)$. The force is purely vertical, balancing gravity; the drag does not contribute. In the string limit ($E=0$), it is not possible to apply a moment, and as pointed out in several previous works Gregory49 , Taberlet19 , Daerr19 , the integral moments about the support due to drag and gravity balance each other; we can also infer that the moment about the center of mass due to drag balances that due to the support force.
It is illustrative to integrate the tangential projection of the linear momentum balance for axially moving equilibria (9), not simplified into (11), but rather rearranged into the form of a pseudomomentum balance MaddocksDichmann94 , OReilly17book , SinghHanna21 (see also Broer70 , Healey96 for strings):
@@ 7 Open questions
While a small amount of bending offers the simplest physically plausible regularization of the vertical singularity, the resulting numerical problem is quite stiff, making comparison with experiment difficult. This suggests the development of a matched asymptotic analysis, likely to be quite different from prior analyses for heavy elastica in which the boundary layer is due to clamping Plunkett67 , StumpvanderHeijden00 , Wolfe93 . We expect the “dolphin nose” Judddolphin , visible in experiments, to emerge at lower bending stiffnesses than those examined here.
@@
The possibility of other forms of solutions, such as undulatory shapes obtained by patching in the moderate drag regime, has not been considered.
@@ 8 Summary and Conclusions
The effects of drag on string shooter solutions can be briefly summarized as follows. For a stationary shape, drag forces do not contribute to the normal force balance, but they change the equilibrium shape and tension. At low values, they do not change the qualitative features of the catenary, in particular its asymptotic approach to verticality, and the blowup of tension of opposite signs on upper and lower branches. Additional physics, such as bending elasticity, is necessary to construct loop solutions. At moderate values, the inertially-shifted tension vanishes along with the curvature at verticality at finite arc length, allowing patching of loop solutions. At high values, the curvature blows up at verticality, with patching still possible.
We have corrected several errors in recent works on axially moving string loops in the presence of gravity. These entail misunderstandings of the role of drag forces, bifurcations, the nature of the vertical singularity, and admissibility of certain solutions. Several example solutions were constructed, either by implementing previously derived analytical results Gregory49 , ChakrabartiHanna16 , or numerically adding bending stiffness. We have also attempted to place the problem in context alongside several other fundamental problems in flexible structure mechanics.
@@ A.1 Derivation of catenary solutions
where $C$ is a constant of integration. For general values of $D$, these expressions are valid as written for the range $\frac{\pi}{2}<\theta<\frac{\pi}{2}$, corresponding in our work to an upper branch of clockwise-travelling string. The compact expressions in the middle are preferred, but the rewritten expressions on the right are intended for further comparison with the work of Gregory Gregory49 , and were obtained using the identities $\sec\theta+\tan\theta=\tan\left(\frac{\pi}{4}+\frac{\theta}{2}\right)$ and $\sec\theta-\tan\theta=\cot\left(\frac{\pi}{4}+\frac{\theta}{2}\right)$.
@@ A.2 Patching of catenary solutions
It turns out that the above procedure for matching positions only works if the upper and lower portions of the curve are the same length, $\ell_{\text{up}}=\ell_{\text{down}}=\tfrac{\ell}{2}=\tfrac{1}{2}$, between the launch point and the vertical singularity. This result is, surprisingly, independent of the launch and return angles, so that, for example, the lengths of any red and green curves in the family shown in [Daerr19, , figure 3] are the same. This equal length condition was found analytically by Gregory Gregory49 . We reproduce an equivalent demonstration in the following subsection.
@@ A.3 Gregory’s demonstration of equal upper and lower lengths
Borrowing the shorthand notation of Gregory49 , we define the quantity appearing in equations (17)-(21) as $\tau\equiv\tan\left(\frac{\pi}{4}+\frac{\theta}{2}\right)$, and note that $\tau(-\tfrac{\pi}{2})=0$. Using (19-21) with a launch angle such that $\frac{\pi}{2}\leq\theta_{1}=\theta_{\text{launch}}<\frac{\pi}{2}$ and $\frac{\pi}{2}\leq\theta_{2}\leq\theta_{\text{launch}}$, we may obtain the shape of the upper portion of the loop ($s_{\text{up}}$, $x_{\text{up}}$, $y_{\text{up}}$) and find $C_{\text{up}}$ by inverting (19) with a chosen $\theta_{\text{launch}}$ and $\ell_{\text{up}}=s(-\frac{\pi}{2})$. Explicitly, we have